OpenStudy (anonymous):

In a rhombus, the difference of the measures of the two angles between a side and the diagonals is 32 degrees, (m<2-m<1=32 degress) What are the measures of the angles of the rhombus?

OpenStudy (campbell_st):

you need to draw a picture are realise the diagonals of a rhombus intersect at right angles. let the angles be a and b  in triangle ADE a + b = 90 ( by angle sum of a triangle since angle AED = 90) and you know the difference of the angles a - b = 32 so you have 2 equations in 2 unknowns that need to be solved a + b = 90 a - b = 32 after solving for a and b the last thing you need to know to find the angles sizes is that the diagonals bisect the angle through which they pass. hope it helps.

OpenStudy (anonymous):

thanks!